y(2x-y+1)dx + x(3x-4y+3)dy=0

1
Expert's answer
2021-11-22T07:40:11-0500

"y(2x-y+1)dx+x(3x-4y+3)dy=0"

"M(x,y)dx+N(x,y)dy=0"

"M=y(2x-y+1)=2xy-y^2+y"

"N=x(3x-4y+3)=3x^2-4xy+3x"

"\frac{dM}{dy}=2x-2y+1"

"\frac{dN}{dx}=6x-4y+3"

"\frac{dM}{dy}\not=\frac{dN}{dx}" ,thus the equation is not exact

"\frac{1}{M}(\frac{dN}{dx}-\frac{dM}{dy})=\frac{6x-4y+3-2x+2y-1}{y(2x-y+1)}"

"=\frac{4x-2y+2}{y(2x-y+1)}"

"=\frac{2(2x-y+1)}{y(2x-y+1)}=\frac{2}{y}" which is a function of y

I.F="e^{\int\frac{2}{y}dy}=e^{2lny}=y^2"

Multiplying both sides of the equation by y2 we get

"y^3(2x-y+1)dx+xy^2(3x-4y+3)dy=0"

"(2xy^3-y^4+y^3)dx+(3x^2y^2-4xy^3+3xy^2)dy=0"

"\int(2xy^3-y^4+y^3)dx=x^2y^3-xy^4+xy^3, taking\ y\ as\ a\ constant"

"\int(3x^2y^2-4xy^3+3xy^2)dy=x^2y^3-xy^4+xy^3, taking\ x\ as\ a\ constant"

Adding the two, the solution of the equation is "x^2y^3-xy^4+xy^3=C"

Where C is an arbitrary constant